OpenStudy (anonymous):

Simplify the given expression: (6 + 2i) − (8 − 3i)

OpenStudy (phi):

the minus sign out front in - (8 -3i) means -1*(8 - 3i) distribute the -1 ... which means multiply each term inside the parens by -1 then combine "like terms" (add up the reals, and add up the imaginaries)

OpenStudy (anonymous):

-8+3i

OpenStudy (phi):

yes, now you have 6 + 2i -8 +3i combine "like terms"

OpenStudy (phi):

in other words, (6-8) + (2i + 3i)

OpenStudy (anonymous):

-2+5i

OpenStudy (phi):

yes

OpenStudy (anonymous):

ok that was given

OpenStudy (anonymous):

wt about this

OpenStudy (anonymous):

Simplify the given expression: (−5 + 3i) • (1 − 2i)

OpenStudy (phi):

do you know FOIL? or in general , how to multiply for example (x+y)(a+b) ?

OpenStudy (anonymous):

-5+6i

OpenStudy (anonymous):

sry that was the wrong question to ask

OpenStudy (anonymous):

-5-6i

OpenStudy (phi):

? what is the question?

OpenStudy (anonymous):

This is it Simplify the given expression: 4 divided by the quantity of 3 minus 2i

OpenStudy (phi):

\[ \frac{4}{3-2i} \] multiply top and bottom by 3+2i can you do that?

OpenStudy (anonymous):

12-8i

OpenStudy (phi):

what did you get for (3-2i)(3+2i) ? (the bottom part)

OpenStudy (phi):

the top part is 4(3+2i) = 12+8i we need the bottom part (3-2i)(3+2i)

OpenStudy (anonymous):

12+8i top, 13 bottom

OpenStudy (phi):

so the answer is \[ \frac{12}{13} + \frac{8}{13} i \] not very pretty or simple, but at least it is in standard a+ bi form

OpenStudy (anonymous):

Next question

OpenStudy (anonymous):

Simplify the given expression: (4 + 2i) − (1 − 7i)

OpenStudy (phi):

same idea as the one at the top. can you do it?

OpenStudy (anonymous):

oh, nvm i got it

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

another one

OpenStudy (anonymous):

(−5 + 3i) • (1 − 2i)

OpenStudy (anonymous):

i messed up on it i think

OpenStudy (phi):

do you know FOIL?

OpenStudy (anonymous):

i got 5i-6

OpenStudy (anonymous):

-5-6i

OpenStudy (phi):

how?

OpenStudy (phi):

you can use the distributive Law or FOIL (same thing, but a way to remember it)

OpenStudy (anonymous):

Are you asking how I got it wrong or how I got to this answer?

OpenStudy (phi):

how did you get that answer? what steps did you do?

OpenStudy (anonymous):

i did the foil method, so i got -6isquared + 5 -5

OpenStudy (anonymous):

but it has to be simplified

OpenStudy (phi):

here is using the distributive law: A (x+y) = Ax + A y (−5 + 3i) • (1 − 2i) in this case, A is (-5 + 3i), and we write (-5 + 3i)* 1 + (-5 + 3i)(-2i) now use the distributive property a second time (twice) the first part 1*(-5 + 3i) is easy : -5 + 3i the 2nd part: -2i*(-5) + (3i)*(-2i) = +10i -6i^2 so we get -5 + 3i +10i -6i^2 which should match the terms you get if you use FOIL next, remember i^2 or i*i is -1 <--- IMPORTANT SIMPLIFICATION in other words, we have -5 + 3i +10i -6*-1

OpenStudy (phi):

now simplify and combine like terms

OpenStudy (anonymous):

oh ok, thx

OpenStudy (anonymous):

what is an extraneous soulution

OpenStudy (anonymous):

oh ok, so here's the question

OpenStudy (anonymous):

4Square root of x minus 3 = 12, solve for x and identify if it is an extraneous solution

OpenStudy (anonymous):

is it not extranuous

OpenStudy (anonymous):

x = 0

OpenStudy (goformit100):

Hello, and A Warm Welcome to Open Study!